# MOSPA: Human Motion Generation Driven by Spatial Audio

### v1

@@ Abstract
Enabling virtual humans to dynamically and realistically respond to diverse auditory stimuli remains a key challenge in character animation, demanding the integration of perceptual modeling and motion synthesis. Despite its significance, this task remains largely unexplored. Most previous works have primarily focused on mapping modalities like speech, audio, and music to generate human motion. As of yet, these models typically overlook the impact of spatial features encoded in spatial audio signals on human motion. To bridge this gap and enable high-quality modeling of human movements in response to spatial audio, we introduce the first comprehensive Spatial Audio-Driven Human Motion (SAM) dataset, which contains diverse and high-quality spatial audio and motion data. For benchmarking, we develop a simple yet effective diffusion-based generative framework for human MOtion generation driven by SPatial Audio, termed MOSPA, which faithfully captures the relationship between body motion and spatial audio through an effective fusion mechanism. Once trained, MOSPA could generate diverse realistic human motions conditioned on varying spatial audio inputs. We perform a thorough investigation of the proposed dataset and conduct extensive experiments for benchmarking, where our method achieves state-of-the-art performance on this task. Our model and dataset will be open-sourced upon acceptance. Please refer to our supplementary video for more details.
[…]
@@ 1 Introduction
To address this overlooked aspect, we propose to model the complex interactions between spatial audio inputs and human motion using a generative model. Since there is no such dataset tailored for this task, we first introduce the SAM dataset (Spatial Audio Motion dataset), which captures diverse human responses to various spatial audio conditions. This dataset is meticulously curated to include a wide range of spatial audio scenarios, enabling the study of motion conditioned on sound field variations. The SAM dataset has a total of more than 9 hours of motion, covering 27 common spatial audio scenarios and more than 70 audio clips. To ensure the diversity of the spatial audio, 480 seconds of motion are captured for each audio clip at different positions in the character space. To ensure diverse motion responses to spatial audio, we introduce 20 distinct motion types (excluding motion genres) and 49 in total when motion genres are included. We visualize samples from SAM in the top row of Fig. 1. See Appendix A for detailed statistics.
We further conduct benchmarking experiments on the proposed dataset, revealing the limitations of existing methods in this setting. To enable spatial audio-driven human motion generation, we introduce MOSPA, a simple yet effective framework tailored for this task. In real-world scenarios, human responses to sound are inherently influenced by spatial perception, intensity variations, directional cues, temporal dynamics, etc. Motivated by this, we generate motion by incorporating features extracted from the input spatial audio signals using mcfee2015librosa . Specifically, to capture intrinsic features across both temporal and spatial dimensions, we mainly utilize Mel-Frequency Cepstral Coefficients (MFCCs)davis1980comparison and Tempogramsgrosche2010cyclic to model the temporal characteristics of the audio. Additionally, we characterize the spatial audio by analyzing the root mean square (RMS) mcfee2015librosa energy, which quantifies signal intensity in audio processing. These features enhance the effective modeling of the spatial and intensity variations of the spatial audio. To capture the distribution of spatial audio features and human motion dynamics effectively, we employ a diffusion-based generative model that ensures strong alignment between the two modalities—human motion and spatial audio signals. Leveraging diffusion models, MOSPA excels at modeling the complex interplay between spatial audio features and human motion. Besides, a residual feature fusion mechanism is employed to model the subtle influences of spatial audio on human movement.
@@ 2 Related Work
Spatial Audio
Many studies have explored spatial audio modeling, perception, and generation zhao2018sound ; gan2019self ; zheng2024bat ; sun2024both ; huang2024modeling ; xu2023sounding . For instance, zhao2018sound utilizes the natural synchronization between visual and audio modalities to learn models that jointly parse sounds and images without manual annotations. gan2019self leverages unlabeled audiovisual data to localize objects, such as moving vehicles, using only stereo sound at inference time. zheng2024bat introduces an approach for reasoning about spatial sounds with large language models. Recently, spatial audio generation has been explored from text sun2024both and video kimvisage . huang2024modeling presents a framework for spatial audio generation, capable of rendering 3D soundfields generated by human actions, including speech, footsteps, and hand-body interactions. xu2023sounding models 3D spatial audio from body motion and speech, while challenges remain to fully capture these dynamics.
Conditional Motion Generation
Extensive efforts have been made into motion synthesis conditioning on user control signals holden2017phase ; starke2022deepphase ; chen2024taming ; wan2023tlcontrol ; zhang2018mode , text zhang2023generating ; zhou2025emdm ; tevet2023human ; guo2022generating ; liao2024rmd ; liu2023plan ; cong2024laserhuman , action xu2023actformer ; guo2020action2motion ; chen2024pay ; huang2024controllable , music tseng2023edge ; zhang2024bidirectional ; le2023controllable ; alexanderson2023listen , speech ao2022rhythmic ; shi2024takes ; huang2024interact ; zhang2024semantic ; ao2023gesturediffuclip ; zhu2023taming , past trajectories feng2024motionwavelet ; chen2023humanmac ; barquero2023belfusion ; yuan2019diverse ; sun2024comusion , etc. We refer readers to  zhu2023human for a detailed survey of motion generation.
@@
While audio signals have been widely used in music- and speech-to-motion tasks, human motion synthesis driven by spatial audio remains largely unexplored. As a result, data-driven methods are highly constrained by limited paired data. The goal of this paper is to develop a comprehensive dataset and a novel approach for high-quality spatial audio-driven motion synthesis.
@@ 3 SAM Dataset
Table 1: Statistics of the SAM dataset. The SAM dataset encompasses 27 common daily spatial audio scenarios, over 20 reaction types excluding the motion genres, and 49 reaction types (see details in Appendix A). The number of subjects covered in SAM is 12, where 5 of them are female and the remaining 7 are male. It is also the first dataset to incorporate spatial audio information, annotated with Sound Source Location (SSL). The total duration of the dataset exceeds 33K seconds.
@@
Dance w/ Melody tang2018dance $\times$ ✓/$\times$ - 21 - 5640
Music2Dance zhuang2022music2dance $\times$ ✓/$\times$ - 55 2 3472
@@
FineDance li2023finedance $\times$ ✓/✓ SMPL-X 52 27 5061K
SAM (Ours) ✓/✓ SMPL-X 55 12 33600
(a) Look for sound source and approach upon hearing miaow at the left hand side.
@@
(d) Walk towards the sound source upon hearing insect noise at the left hand side.
@@
(f) Look for sound source upon hearing the phone ring at the left hand side.
@@
We utilize a Vicon motion capture system Vicon to collect motion data and spatial audio signals. The motion capture is performed in a space of $5\text{m}\times 10\text{m}\times 3\text{m}$, with 28 cameras recording at a frame rate of 120 Hz; See Fig. 4. In SAM, each audio clip is paired with 16 randomly sampled relative locations of the sound source. For each location, we capture three motion sequences corresponding to different reaction intensities: dull, neutral, and sensitive, resulting in a total of 48 motion sequences, each lasting 10 seconds.
@@
Data Processing.
All motion and audio clips are precisely aligned. The motions are re-targeted and converted from the original BVH format in Vicon to the SMPL-X pavlakos2019expressive format. SMPL-X is a parametric 3D human body model that encompasses the body, hands, and face, comprising $N=10,475$ vertices and $K=55$ joints. Given shape parameters $\beta$ and pose parameters $\theta$, the SMPL-X model generates the corresponding body shape and pose through forward dynamics. We extract the locations of the sound sources in each motion clip. The sound source locations are then transformed into the local space of the character aligned with the SMPL-X local coordinate system (also known as the local frame).
@@ 4 Method
We introduce MOSPA, a diffusion-based probabilistic model that serves as a baseline for this novel task of spatial audio-driven human motion generation. First, we extract spatial audio features $\mathbf{a}$ using a feature extractor mcfee2015librosa . During motion generation, the extracted spatial audio feature $\mathbf{a}$ is combined with the sound source location $\mathbf{s}$ and the motion genre $g$ as conditioning inputs. These inputs are passed to a denoiser $\mathcal{G}$, which is trained to reconstruct the original clean motion vector $\mathbf{\hat{x_{0}}}$ by denoising the given noisy motion vector $\mathbf{x_{t}}$ at time step $t$. Mathematically, we have $\mathbf{\hat{x_{0}}}=\mathcal{G}(\mathbf{x_{t}},t;\mathbf{a},\mathbf{s},g)$
@@ 4.1 Feature Representation
The two vectors that we care the most about are the audio feature vector $\mathbf{a}$ and the motion vector $\mathbf{x}$. We carefully design the structure of the two vectors in MOSPA.
Spatial Audio Feature Extraction. We first extract a range of audio features that capture intensity, temporal dynamics, and spatial characteristics. Inspired by siyao2022bailando , our feature set primarily includes Mel-frequency cepstral coefficients (MFCC), MFCC delta, constant-Q chromagram, short-time Fourier transform (STFT) of the chromagram, onset strength, tempogram, and beats davis1980comparison ; schorkhuber2010constant ; grosche2010cyclic ; mcfee2015librosa ; bock2013maximum ; ellis2007beat . On top of these audio features, we additionally add the root mean square (RMS) energy $E_{rms}$ of the audio mcfee2015librosa , and the active frames $F_{active}$ defined as $F_{active}=E_{rms}>0.01$ to capture the distance information of the audio. The dimension of the audio feature vector for each ear is 1136. By concatenating the features from both ears, we obtain a combined feature vector $\mathbf{a}$ of dimension 2272. The detailed construction of the audio vector can be viewed in Appendix B.1.
@@ 4.2 Framework
where $\alpha_{t}\in(0,1)$. The model then learns to gradually denoise a noisy motion vector $\mathbf{x_{t}}$ in $t$ time steps, i.e. $p(\mathbf{x_{t-1}}|\mathbf{x_{t}})$. We directly predict the clean sample $\mathbf{\hat{x_{0}}}$ in each diffusion step $\mathbf{\hat{x_{0}}}=\mathcal{G}(\mathbf{x_{t}},t;\mathbf{a},\mathbf{s},g)$, where $\mathbf{a}$ is the audio features, $\mathbf{a}$ is the sound source location and $g$ is the motion genre. This strategy, employed by  tevet2023human ; chen2024taming ; ramesh2022hierarchical ; xie2023omnicontrol , has been proved to be more efficient and accurate than predicting the noise $\epsilon_{t}$, suggested by  ho2020denoising . We employ an encoder-only transformer to reverse the diffusion process and predict the clean samples. The timestep, motions and conditions are projected to the same latent dimension by separate feed-forward networks, and then concatenated together to provide the full tokens $\mathbf{z}$. The tokens are positionally embedded afterward and input into a transformer to get the output $\mathbf{\hat{z}}$. The predicted clean sample is thus extracted from the last $T$ tokens of $\mathbf{\hat{z}}$ by inputting it to another feed-forward network, where $T=240$ is the length of the motion; see Fig. 5.
@@ 4.4 Implementation Details
[…]
5 Experiments
5.1 Comparisons
@@
[…]
@@
[…]
@@
Quantitative Results. The quantitative results are reported in Tab. 2. MOSPA achieves the best performance as shown by the lowest FID value and the highest R-precision values. Also, our generated motions exhibit the closest diversity and APD dou2023c values compared with the Real Motion, demonstrating the effectively balanced variation and precision. Bailando siyao2022bailando has the worst performance among the four baselines in practice, as illustrated by the extremely high FID. The model possibly lacks an ability to perceive commonly heard sounds other than music and also spatial information of the audio. Our method, overall speaking, still demonstrates competitive performance in spatial audio conditioned motion generation, which is proved by the low values in precision-related metrics.
@@
User Study. We conducted a user study with 25 participants to assess the perceptual quality of motion generation. Participants evaluated five models (MOSPA, EDGE, POPDG, LODGE, Bailando) alongside ground truth (GT), selecting the best motion for: 1) Human Intent Alignment: Does the motion align with real-world intent? 2) Motion Quality: Which has the highest movement quality? 3) GT Similarity:Which best matches the GT motion? We provided GT motion and a textual description for reference. As shown in Fig. 7, MOSPA (Ours) outperforms all baselines across all criteria, while LODGE and Bailando received the fewest selections, indicating limitations in generating realistic, semantically meaningful motions. See more details in Appendix C.

### v2

@@ Abstract
Enabling virtual humans to dynamically and realistically respond to diverse auditory stimuli remains a key challenge in character animation, demanding the integration of perceptual modeling and motion synthesis. Despite its significance, this task remains largely unexplored. Most previous works have primarily focused on mapping modalities like speech, audio, and music to generate human motion. As of yet, these models typically overlook the impact of spatial features encoded in spatial audio signals on human motion. To bridge this gap and enable high-quality modeling of human movements in response to spatial audio, we introduce the first comprehensive Spatial Audio-Driven Human Motion (SAM) dataset, which contains diverse and high-quality spatial audio and motion data. For benchmarking, we develop a simple yet effective diffusion-based generative framework for human MOtion generation driven by SPatial Audio, termed MOSPA, which faithfully captures the relationship between body motion and spatial audio through an effective fusion mechanism. Once trained, MOSPA can generate diverse realistic human motions conditioned on varying spatial audio inputs. We perform a thorough investigation of the proposed dataset and conduct extensive experiments for benchmarking, where our method achieves state-of-the-art performance on this task. Our code and model are publicly available at our website.
footnotetext: $\dagger$} denote equal contributions and corresponding authors.
@@ 1 Introduction
To address this overlooked aspect, we propose to model the complex interactions between spatial audio inputs and human motion using a generative model. Since there is no such dataset tailored for this task, we first introduce the SAM dataset (Spatial Audio Motion dataset), which captures diverse human responses to various spatial audio conditions. This dataset is meticulously curated to include a wide range of spatial audio scenarios, enabling the study of motion conditioned on sound field variations. The SAM dataset has a total of more than 9 hours of motion, covering 27 common spatial audio scenarios and more than 70 audio clips. To ensure the diversity of the spatial audio, around 480 seconds of motion were captured for each audio clip at different positions in the character space. To ensure diverse motion responses to spatial audio, we introduce 20 distinct motion types (excluding motion genres) and 49 in total when including motion genres. We visualize samples from SAM in the top row of Fig. 1. See Appendix A for detailed statistics.
We further conduct benchmarking experiments on the proposed dataset, revealing the limitations of existing methods in this setting. To enable spatial audio-driven human motion generation, we introduce MOSPA, a simple yet effective framework tailored for this task. In real-world scenarios, human responses to sound are inherently influenced by spatial perception, intensity variations, directional cues, temporal dynamics, etc. Motivated by this, we generate motion by incorporating features extracted from the input spatial audio signals using mcfee2015librosa . Specifically, to capture intrinsic features across both temporal and spatial dimensions, we mainly utilize Mel-Frequency Cepstral Coefficients (MFCCs)davis1980comparison and Tempogramsgrosche2010cyclic to model the temporal characteristics of the audio. Additionally, we characterize the spatial audio by analyzing the root mean square (RMS) mcfee2015librosa energy, which quantifies signal intensity in audio processing.
These features enhance the effective modeling of the spatial and intensity variations of the spatial audio. To capture the distribution of spatial audio features and human motion dynamics effectively, we employ a diffusion-based generative model that ensures strong alignment between the two modalities—human motion and spatial audio signals. Leveraging diffusion models, MOSPA excels at modeling the complex interplay between spatial audio features and human motion. Besides, a residual feature fusion mechanism is employed to model the subtle influences of spatial audio on human movement.
@@ 2 Related Work
Spatial Audio. Many studies have explored spatial audio modeling zhao2018sound ; gan2019self ; zheng2024bat ; sun2024both ; huang2024modeling ; xu2023sounding ; lan2024acoustic ; lan2025resoundingacousticfieldsreciprocity . For instance, zhao2018sound utilizes the natural synchronization between visual and audio modalities to learn models that jointly parse sounds and images without manual annotations. gan2019self leverages unlabeled audiovisual data to localize objects, such as moving vehicles, using only stereo sound at inference time. zheng2024bat reason about spatial sounds with large language models. Recently, spatial audio generation has been explored from text sun2024both and video kim2025visage . xu2023sounding propose a method to model 3D spatial audio from body motion and speech. huang2024modeling presents a framework for spatial audio generation, capable of rendering 3D soundfields generated by human actions, including speech, footsteps, and hand-body interactions. Despite progress on spatial-audio tasks, generating human motion from spatial audio remains largely underexplored.
Conditional Motion Generation. Extensive efforts have been made into motion synthesis conditioning on user control signals holden2017phase ; starke2022deepphase ; chen2024taming ; wan2023tlcontrol ; zhang2018mode , text zhang2023generating ; zhou2025emdm ; tevet2023human ; guo2022generating ; liao2024rmd ; liu2023plan ; cong2024laserhuman , action xu2023actformer ; guo2020action2motion ; chen2024pay ; huang2024controllable , music tseng2023edge ; zhang2024bidirectional ; le2023controllable ; alexanderson2023listen , speech ao2022rhythmic ; shi2024takes ; huang2024interact ; zhang2024semantic ; ao2023gesturediffuclip ; zhu2023taming , past trajectories feng2024motionwavelet ; chen2023humanmac ; barquero2023belfusion ; yuan2019diverse ; sun2024comusion , etc. We refer readers to  zhu2023human for a detailed survey of motion generation.
@@
While audio cues power many music- and speech-to-motion methods, the use of spatial audio for human motion synthesis is still scarcely studied. As a result, data-driven methods are highly constrained by limited paired data. The goal of this paper is to develop a comprehensive dataset and a novel approach for high-quality spatial audio-driven motion synthesis.
@@ 3 SAM Dataset
Table 1: Statistics of the SAM dataset. The SAM dataset encompasses 27 common daily spatial audio scenarios, over 20 reaction types excluding the motion genres, and 49 reaction types (see details in Appendix A). The number of subjects covered in SAM is 12, where 5 of them are female and the remaining 7 are male. It is also the first dataset to incorporate spatial audio information, annotated with Sound Source Location (SSL). The total duration of the dataset exceeds 34K seconds.
@@
Dance with Melody tang2018dance $\times$ $\times -$ 21 - 5640
DanceNet zhuang2022music2dance $\times$ $\times -$ 55 2 3472
@@
FineDance li2023finedance $\times$ ✓/✓ SMPL & hand joints 52 27 52560
SAM (Ours) ✓/✓ SMPL-X 55 12 34356
(a) Look for sound source and approach upon hearing miaow at the left-hand side.
@@
(d) Walk towards the sound source upon hearing insect noise at the left-hand side.
@@
(f) Look for sound source upon hearing the phone ring at the left-hand side.
@@
Data Capture Settings. We utilize a Vicon motion capture system Vicon to collect motion data and spatial audio signals. The motion capture is performed in a semi-open cage having a space of approximately $5\text{m}\times 10\text{m}\times 3\text{m}$, a structure covered by rope nets, with 28 mocap cameras mounted on the ropes and vertical supports recording at a frame rate of 120 Hz; See Fig. 4. The surrounding walls are standard painted concrete, resulting in a setting that resembles a typical indoor environment. In SAM, each audio clip is associated with 16 randomly sampled relative sound source locations, defined by combinations of different speakers and spatial positions relative to the subject. For each location, we capture three motion sequences corresponding to different reaction intensities: dull, neutral, and sensitive, resulting in a total of 48 motion sequences, each lasting 10 seconds.
@@
Data Processing. All motion and audio clips are precisely aligned. The motions are re-targeted and converted from the original BVH format in Vicon to the SMPL-X pavlakos2019expressive format. SMPL-X is a parametric 3D human body model that encompasses the body, hands, and face, comprising $N=10,475$ vertices and $K=55$ joints. Given shape parameters $\beta$ and pose parameters $\theta$, the SMPL-X model generates the corresponding body shape and pose through forward dynamics. We extract the locations of the sound sources in each motion clip. The sound source locations are then transformed into the local space of the character aligned with the SMPL-X local coordinate system (a.k.a the local frame).
@@ 4 Method
We introduce MOSPA, a diffusion-based probabilistic model that serves as a baseline for this novel task of spatial audio-driven human motion generation. First, we extract spatial audio features $\mathbf{a}$ using a feature extractor mcfee2015librosa . During motion generation, the extracted spatial audio feature $\mathbf{a}$ is combined with the sound source location $\mathbf{s}$ and the motion genre $g$ as conditioning inputs. These inputs are passed to a denoiser $\mathcal{G}$, which is trained to reconstruct the original clean motion vector $\mathbf{\hat{x_{0}}}$ by denoising the given noisy motion vector $\mathbf{x_{t}}$ at time step $t$. Mathematically, we have $\mathbf{\hat{x_{0}}}=\mathcal{G}(\mathbf{x_{t}},t;\mathbf{a},\mathbf{s},g)$.
@@ 4.1 Feature Representation
The two key vectors are the audio feature vector $\mathbf{a}$ and the motion vector $\mathbf{p}$. We carefully designed the structure of the two vectors in MOSPA.
Spatial Audio Feature Extraction. We first extract a range of audio features that capture intensity, temporal dynamics, and spatial characteristics. Inspired by siyao2022bailando , our feature set primarily includes Mel-frequency cepstral coefficients (MFCCs), MFCC delta, constant-Q chromagram, short-time Fourier transform (STFT) of the chromagram, onset strength, tempogram, and beats davis1980comparison ; schorkhuber2010constant ; grosche2010cyclic ; mcfee2015librosa ; bock2013maximum ; ellis2007beat . On top of these audio features, we additionally add the root mean square (RMS) energy $E_{rms}$ of the audio mcfee2015librosa , and the active frames $F_{active}$ defined as $F_{active}=E_{rms}>0.01$ to capture the distance information of the audio. The dimension of the audio feature vector for each ear is 1136. By concatenating the features from both ears, we obtain a combined feature vector $\mathbf{a}$ of dimension 2272. The detailed construction of the audio vector can be viewed in Appendix B.1.
@@ 4.2 Framework
where $\alpha_{t}\in(0,1)$. The model then learns to gradually denoise a noisy motion vector $\mathbf{x_{t}}$ in $t$ time steps, i.e. $p(\mathbf{x_{t-1}}|\mathbf{x_{t}})$. We directly predict the clean sample $\mathbf{\hat{x_{0}}}$ in each diffusion step $\mathbf{\hat{x_{0}}}=\mathcal{G}(\mathbf{x_{t}},t;\mathbf{a},\mathbf{s},g)$, where $\mathbf{a}$ is the audio features, $\mathbf{s}$ is the sound source location and $g$ is the motion genre. This strategy, employed by  tevet2023human ; chen2024taming ; ramesh2022hierarchical ; xie2023omnicontrol , has been proved to be more efficient and accurate than predicting the noise $\epsilon_{t}$, suggested by  ho2020denoising . We employ an encoder-only Transformer to reverse the diffusion process and predict the clean samples. The timestep, motion, and conditioning signals are each projected into the same latent dimension using separate feed-forward networks. Random masks are applied to the audio features $\mathbf{a}$ and the sound source location (SSL) $\mathbf{s}$, after which all components are concatenated to form the complete token sequence $\mathbf{z}$. The tokens are positionally embedded afterward and input into a transformer to get the output $\mathbf{\hat{z}}$. The predicted clean sample is thus extracted from the last $T$ tokens of $\mathbf{\hat{z}}$ by inputting it to another feed-forward network, where $T=240$ is the length of the motion; see Fig. 5.
@@ 4.4 Implementation Details
In MOSPA, the diffusion model is a transformer-based diffusion network chen2024taming ; tevet2023human ; zhou2025emdm . The encoder transformer is configured with a latent dimension of 512, 8 heads, and 4 layers. We employ AdamW loshchilov2017decoupled as the optimizer with an initial value of $1\text{\times}{10}^{-4}$. The number of denoising steps used is 1000, and the noise schedule is cosine. The training phase concludes after $6,000$ epochs. Exceeding these recommended epoch counts may degrade model quality due to overfitting. The entire training process requires approximately 18 hours on a single RTX 4090 GPU with a batch size of 128.
@@
Method $\text{R-precision}\uparrow$ $\text{FID}\downarrow$ $\text{Diversity}\rightarrow$ $\text{APD}\rightarrow$
@@
Latent Dim Head Num Diff Steps Genre $\text{R-precision}\uparrow$ $\text{FID}\downarrow$ $\text{Diversity}\rightarrow$ $\text{APD}\rightarrow$
@@
5 Experiments
5.1 Comparisons
Quantitative Results. The quantitative results are reported in Tab. 2. MOSPA achieves the best performance as shown by the lowest FID value and the highest R-precision values. Also, our generated motions exhibit the closest diversity and APD dou2023c values compared with the Real Motion, demonstrating the effectively balanced variation and precision. Bailando siyao2022bailando has the worst performance among the four baselines in practice, as illustrated by the extremely high FID. The model possibly lacks the ability to perceive commonly heard sounds other than music and also the spatial information of the audio. Our method, overall speaking, still demonstrates competitive performance in spatial audio conditioned motion generation, which is proved by the low values in precision-related metrics and the high values in diversity-related metrics.
@@ 5.1 Comparisons
User Study. We conducted a user study with 25 participants to assess the perceptual quality of motion generation. Participants evaluated five models (MOSPA, EDGE, POPDG, LODGE, Bailando) alongside ground truth (GT), selecting the best motion for: 1) Human Intent Alignment: Does the motion align with real-world intent? 2) Motion Quality: Which has the highest movement quality? 3) GT Similarity:Which best matches the GT motion? We provided GT motion and a textual description for reference. As shown in Fig. 7, MOSPA outperforms all baselines across all criteria, while LODGE and Bailando received the fewest selections, indicating limitations in generating realistic, semantically meaningful motions. See more details in Appendix C.